In order ti start a small business, a student takes out a simple interest loan for $4000 for 9 months at a rate of 8.25%.
a-) How much interest must the student pay?
b-) Find the future value of the loan.

Answer by jim_thompson5910(35256)

FV = PV*(1+rt)

FV = 4000*(1+0.0825*(9/12))

FV = 4247.5

a)
The student must pay 4247.5 - 4000 = 247.50 dollars in interest

b)
The future value of the loan is 4247.50 dollars.